OpenStudy (anonymous):

How do you solve solve cos2x+cos4x=0 ?

OpenStudy (anonymous):

Let cos²x = u. You get u+u² = 0 Factor u out: u(u+1) = 0 u = 0, -1 Now you have cos²x = 0 and cos²x = -1 Does this help you?

OpenStudy (anonymous):

I thought you were supposed to use the double-angle properties?

OpenStudy (anonymous):

Oh, I though you mean cos²x, sorry. I know you actually mean cos2x. Yes, go ahead and use double-angle properties.

OpenStudy (anonymous):

can you explain that? that's what I'm having trouble with

OpenStudy (anonymous):

like the first step to solve

OpenStudy (anonymous):

We know that \(\cos(\alpha + \beta) = \cos\alpha \cos\beta-\sin\alpha\sin\beta\) So cos2x = cos²x - sin²x cos2x + cos4x = (cos²x - sin²x) + (cos²2x - sin²2x) = (cos²x - sin²x) + ((cos²x - sin²x)² - (2cosxsinx)²) I honestly don't know if this is correct path to solution, though. Hope I helped.

OpenStudy (anonymous):

Also \(\sin(\alpha + \beta) = \cos\alpha \sin\beta + \cos\beta\sin\alpha\) So sin2x = cosxsinx + cosxsinx = 2cosxsinx

OpenStudy (anonymous):

thank you for your input